8.   In a population of 12,000 individuals, 300 men are afflicted with a recessive, X-linked disease. How many woman would be expected to be afflicted in this population? (Assume the population is 50% men and 50% women.)

a.   30

b.   7.5

c.   60

d.   15

e.   3.75
Answer #1

The nunber of male individuals in a population are 12,000/2=6,000

The male has XY chromosome.

300 male are affected out of 6,000 individual.

Genotypic or allelic frequency of disease allele is same in the men population = 300/6,000=0.05

The female individual afflicted with this disease are (0.05)^2×6,000= 15

Right answer is d)15